Evaluate this integral x(x²+4)dx / x⁴+9

1
Expert's answer
2022-02-22T05:21:55-0500
"\int \dfrac{x(x^2+4)}{x^4+9}dx"

Substitute "x^2=u, 2xdx=du"


"\int \dfrac{x(x^2+4)}{x^4+9}dx=\dfrac{1}{2}\int \dfrac{u+4}{u^2+9}du"

"=\dfrac{1}{2}\int \dfrac{u}{u^2+9}du+2\int \dfrac{du}{u^2+9}"

"\dfrac{1}{2}\int \dfrac{u}{u^2+9}du"

"t=u^2+9, dt=2udu"

"\dfrac{1}{2}\int \dfrac{u}{u^2+9}du=\dfrac{1}{4}\int \dfrac{dt}{t}="

"=\dfrac{1}{4}\ln |t|+C_1=\dfrac{1}{4}\ln(u^2+9)+C_1"

"=\dfrac{1}{4}\ln(x^4+9)+C_1"

"2\int \dfrac{du}{u^2+9}"

"s=\dfrac{u}{3}, ds=\dfrac{1}{3}du"

"2\int \dfrac{du}{u^2+9}=2\int \dfrac{3ds}{9s^2+9}=\dfrac{2}{3}\int \dfrac{ds}{1+s^2}"

"=\dfrac{2}{3}\tan^{-1}(s)+C_2=\dfrac{2}{3}\tan^{-1}(\dfrac{u}{3})+C_2"


"=\dfrac{2}{3}\tan^{-1}(\dfrac{x^2}{3})+C_2"

Then


"\int \dfrac{x(x^2+4)}{x^4+9}dx=\dfrac{1}{4}\ln(x^4+9)+\dfrac{2}{3}\tan^{-1}(\dfrac{x^2}{3})+C"